40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Apr 26 22:56:01 2011, kiwamu, Update, LSC, c1iscaux2 and c1iscaux restrated 
    Reply  Wed Apr 27 21:19:25 2011, kiwamu, Update, LSC, LSC whitening for PD1-4 Screenshot-1.png
Message ID: 4570     Entry time: Tue Apr 26 22:56:01 2011     Reply to this: 4577
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: c1iscaux2 and c1iscaux restrated 

While checking whitening filters on the LSC rack, I found some epics controls for the whitening looked not working.

So I powered two crates off : the top one and the bottom one on 1Y3 rack.

These crates contain c1iscaux and c1iscaux2. Then powered them on. But it didn't solve the issue.

ELOG V3.1.3-